OpenStudy (anonymous):

what is the anwser to [(21+6)-32)]/9*2

OpenStudy (akashdeepdeb):

-10/9. If the 2 is multiplied to the numerator! :)

OpenStudy (nurali):

\[\frac{ (21+6)-32 }{ 9*2 }\] \[\frac{ 27-32 }{ 18 }\] \[\frac{ -5 }{ 18 }\]

OpenStudy (akashdeepdeb):

And if the 2 is multiplied to the denominator then the answer is -5/18 :)

OpenStudy (anonymous):

Solve the inner brackets first (21+6) =27 27-32= -5 so -5/ (9*2) = -5/18

OpenStudy (anonymous):

-5 / 18 =

OpenStudy (akashdeepdeb):

Kudos?